(996) Problem 1: To start a car engine, the car battery moves 3.95 x 1021 electrons through the starter motor. tatus for ew atus theesperta.com-tracking id7072-FE-99-48-9962-19549 In accordance with Expert TAs Terms of Service copying this information to any solutions sharing website is strictly forbidden Doing so may result in termination of your Expet TA Account How many coulombs of charge are moved? Grade Summar Deductions Potential tan) Submissions Attempts remain sino cos0 sin 41 5 6
Answer #1

Solution) Coulombs of charge moved, Q = no. of electrons \times charge of electron = 3.95 \times 10^21 \times 1.6 \times 10^(-19)

= 632 C.